Percent of Sales Method
At the end of the current year, Accounts Receivable has a balance of $3,750,000, Allowance for Doubtful Accounts has a credit balance of $22,750, and sales for the year total $48,400,000. Bad Debt Expense is estimated at ¾ of 1% of sales.
a. Determine the amount of the adjusting entry for uncollectible accounts.
b. Determine the adjusted balances of Accounts Receivable, Allowance for Doubtful Accounts, and Bad Debt Expense.
Adjusted Balance
Debit (Credit)
Accounts Receivable $
Allowance for Doubtful Accounts $
Bad Debt Expense $
c. Determine the net realizable value of accounts receivable.
$
Answer #1

a)

48,400,000 * 1% * 3/4 = 363,000

b)

Accounts Receivable = 3,750,000

Allowance for Doubtful Accounts = 22750+363,000

                                                       = 385,750

Bad debts expense = 363,000

c) Net realizable value = accounts receivable - Allowance for Doubtful Accounts

                                     = 3,750,000 - 385,750

                                     = 3,364,250

I HOPE IT USEFUL TO YOU IF YOU HAVE ANY DOUBT PLZ COMMENT GIVE ME UP-THUMB. THANKS..